# Measurement-free code-switching for low overhead quantum computation using permutation invariant codes

### v2

@@ Abstract
Transversal gates on quantum error correction codes have been a promising approach for fault-tolerant quantum computing, but are limited by the Eastin-Knill no-go theorem. Existing solutions like gate teleportation and magic state distillation are resource-intensive. We present a measurement-free code-switching protocol for universal quantum computation, switching between a stabiliser code for transversal Cliffords and a permutation-invariant code for transversal non-Cliffords that are logical $Z$ rotations for any rational multiple of $\pi$. The novel non-Clifford gates enabled by this code-switching protocol enable implementation of a universal gate set more efficient than the Clifford$+T$ gate set. To achieve this, we present a protocol for performing controlled-NOTs between the codes using near-term quantum control operations that employ a catalytic bosonic mode.
@@ V Implementing entangling logical gates
The different GPG’s discussed above enable use of different logical operations for code switching. The non-linear GPG can implement logical CNOT gates between a non even-odd PI code, like the PI-7 code, and an even-odd stabilizer code (see Appendix E). However, the dispersive interactions between the spins and the cavity mode necessary to instantiate the non-linear GPG are not directly accessible in many physical systems [39], and the dispersive interaction strengths tend to be smaller than linear interaction strengths implying slower gates. We therefore focus most of our analysis on using linear GPGs.
@@ V.1.2 Implementing the logical $\overline{\rm{H}}$ gate
where $\mathcal{E}_{\rm ph}(\rho)=F_{\rm ph}C_{N-1}(Z)\rho C^{\dagger}_{N-1}(Z)$ is the erroneous map of the middle phase gate in Eq. (23). $\mathcal{E}_{\rm l-GPG}$ ($\mathcal{E}^{\rm R}_{\rm l-GPG}$) is the (reverse of) preparation mapping associated with $W$ ($W^{\dagger}$) in the presence of losses, as detailed in Eq. (18). Here, for simplicity, we optimize parameters solely for the preparation process and apply the same parameters to the reverse mapping without additional optimization. The fidelity measure $F_{\rm pro}(\mathcal{E}_{\overline{H}},\overline{\mathcal{H}})$ is computed based on process matrices, which necessitates the superoperator formalism for each mapping in Eq. (24). The detailed calculations are provided in Appendix D.
In Fig. 2, we plot the infidelity for: the implementation error of the Hadamard gate (orange), and the preparation error (blue) for the PI-11 code as a function of the cooperativity $C$. The simulation is performed in the presence of loss errors as described in Eq. (16). Following the approximation in Ref. [55] (Eq. (25)), the infidelity of the $\overline{H}$ implementation can be estimated as $1-F_{\rm pro}(\mathcal{E}_{\overline{H}},\overline{\mathcal{H}})\sim 21.78\times N/\sqrt{C}$, where we consider $18$ non-trivial GPG operations for the preparation step and its inverse, with $\theta=\pi/2$. The infidelity for the middle $C_{N-1}(Z)$ gate is approximated by $1.8\times N/\sqrt{C}$. Similarly, the infidelity for state preparation is estimated by $1-\langle+_{\rm pi11}|\rho|+_{\rm pi11}\rangle\sim 9.99\times N/\sqrt{C}$, where $9$ non-trivial GPG gates are considered, again with $\theta=\pi/2$. Our simulation results indicate a somewhat better performance with the infidelity scaling as $1-F_{\rm pro}(\mathcal{E}_{\overline{H}},\overline{\mathcal{H}})\sim 8.29\times N/C^{0.4985}$ and $1-\langle+_{\rm pi11}|\rho|+_{\rm pi11}\rangle\sim 2.80\times N/C^{0.4953}$, indicating lower infidelity compared to the approximation.
@@ VI Approximation of the super golden gate
The most efficient single-qubit gate set is given by representations of the binary icosahedral group $2I$ (which can be implemented with the Pollatsek-Ruskai code) and the super golden gate $\tau_{60}$ given in (6). Here we give an approximate decomposition of $\tau_{60}$ which can be achieved by code-switching between bg-PI codes, the Pollatsek-Ruskai code that implements 2I, and PI codes that implement the transversal $T$ gate.
@@
VII Conclusion
@@
We have presented a method for code switching between a stabilizer code and a permutation invariant code, and along the way have introduced a family of PI codes with a tunable set of transversal, non-Clifford gates. It is noteworthy that our measurement-free code-switching scheme is also applicable to transitions between two stabiliser codes, provided they satisfy the even-odd quantum code structure. In comparison to the measurement-based approach, our scheme eliminates the overhead associated with measurements and reduces the gate operation complexity by utilising collective gate implementations. We leave an analysis of to what extent our method can be made fault-tolerant to future work.
VIII Acknowledgements
@@
Y.O. acknowledges support from EPSRC (Grant No. EP/W028115/1). G.K.B. and Y.J. acknowledge support from the Australian Research Council Centre of Excellence for Engineered Quantum Systems (Grant No. CE 170100009).
@@ Appendix A AAB+ code
$T_{b}^{\otimes n}|1_{b,g}\rangle$ $=e^{ig\pi/b}|0_{b,g}\rangle.$ (32)
@@ Appendix C Transversal non-Clifford logical operations
Appendix D Fidelity measure for Hadamard gate process

### v3

@@ Abstract
Transversal gates on quantum error correction codes have been a promising approach for fault-tolerant quantum computing, but are limited by the Eastin-Knill no-go theorem. Existing solutions like gate teleportation and magic state distillation are resource-intensive. We present a measurement-free code-switching protocol for universal quantum computation, switching between a stabiliser code for transversal Cliffords and a permutation-invariant code for transversal non-Cliffords that are logical $Z$ rotations for any rational multiple of $\pi$. The novel non-Clifford gates enabled by this code-switching protocol enable implementation of a universal gate set more efficient than the Clifford$+T$ gate set in terms of gate-count. To achieve this, we present a protocol for performing controlled-NOTs between the codes using near-term quantum control operations that employ a catalytic bosonic mode.
@@ II.1 Transversal gates on PI codes
For higher distance codes that support logical $Z(\pi g/b)$ rotations with transversal gates, we introduce $(b,g,m)$-PI codes, which generalize $(b,g)$-PI codes. These codes use $N=2bm+g$ qubits. In general, for any positive integer $m$, we define
$|0_{b,g,m}\rangle=\sum_{k=0}^{m}\sqrt{\binom{m}{k}}\frac{\gamma_{b,g,m,k}}{2^{m}\sqrt{(2m-1)!!}}|D^{N}_{2kb}\rangle,$ (8)
and correspondingly $|1_{b,g,m}\rangle=X^{\otimes N}|0_{b,g,m}\rangle$, where for $k=0,\dots,m$, we define
$\gamma_{b,g,m,k}$ $=b^{-m/2}\prod_{i=k+1}^{m}\sqrt{2ib-g}\prod_{j=m-k+1}^{m}\sqrt{2jb+g}.$ (9)
Note that the $(b,g,1)$-PI codes is a $(b,g)$-PI codes. We prove in the Appendix that when $g,2b-g\geq 2t+1$ and $m\geq t$, a $(b,g,m)$-PI code has distance at least $12t+$ 1, and hence corrects $t$ errors. These $(b,g,m)$-PI codes are equivalent to the code of Aydin et al. only when $m=1$. For $m\geq 2$, the $(b,g,m)$-PI codes form a new family of PI codes.
For $m=2$, we find generalized bg codes of the form
$|0_{b,g,2}\rangle=$ $\frac{\sqrt{(2b-g)(4b-g)}}{4b\sqrt{3}}|D^{N}_{0}\rangle$
$+\frac{\sqrt{(4b+g)(4b-g)}}{2b\sqrt{6}}|D^{N}_{2b}\rangle$
$+\frac{\sqrt{(2b+g)(4b+g)}}{4b\sqrt{3}}|D^{N}_{4b}\rangle.$ (10)
For $m=3$, we find generalized bg codes of the form
$|0_{b,g,3}\rangle=$ $\frac{\sqrt{(2b-g)(4b-g)(6b-g)}}{8b\sqrt{15b}}|D^{N}_{0}\rangle$
$+\frac{\sqrt{(4b-g)(6b+g)(6b-g)}}{8b\sqrt{5b}}|D^{N}_{2b}\rangle$
$+\frac{\sqrt{(4b+g)(6b+g)(6b-g)}}{8b\sqrt{5b}}|D^{N}_{4b}\rangle$
$+\frac{\sqrt{(2b+g)(4b+g)(6b+g)}}{8b\sqrt{15b}}|D^{N}_{6b}\rangle.$ (11)
For $m=4$, we find
$|0_{b,g,4}\rangle=$ $\frac{\sqrt{(2b-g)(4b-g)(6b-g)(8b-g)}}{16b^{2}\sqrt{105}}|D^{N}_{0}\rangle$
$+\frac{\sqrt{(4b-g)(6b-g)(8b+g)(8b-g)}}{8b^{2}\sqrt{105}}|D^{N}_{2b}\rangle$
$+\frac{\sqrt{(6b+g)(6b-g)(8b+g)(8b-g)}}{8b\sqrt{70}}|D^{N}_{4b}\rangle$
$+\frac{\sqrt{(4b+g)(6b+g)(8b+g)(8b-g)}}{8b\sqrt{105}}|D^{N}_{6b}\rangle$
$+\frac{\sqrt{(2b+g)(4b+g)(6b+g)(8b+g)}}{16b\sqrt{105}}|D^{N}_{8b}\rangle,$ (12)
and $|1_{b,g,4}\rangle=X^{\otimes N}|0_{b,g,4}\rangle$.
@@ V Implementing entangling logical gates
The different GPG’s discussed above enable use of different logical operations for code switching. The non-linear GPG can implement logical CNOT gates between a non even-odd PI code, like the PI-7 code, and an even-odd stabilizer code (see Appendix G). However, the dispersive interactions between the spins and the cavity mode necessary to instantiate the non-linear GPG are not directly accessible in many physical systems [39], and the dispersive interaction strengths tend to be smaller than linear interaction strengths implying slower gates. We therefore focus most of our analysis on using linear GPGs.
@@ V.1.2 Implementing the logical $\overline{\rm{H}}$ gate
where $\mathcal{E}_{\rm ph}(\rho)=F_{\rm ph}C_{N-1}(Z)\rho C^{\dagger}_{N-1}(Z)$ is the erroneous map of the middle phase gate in Eq. (28). Here, $F_{\rm ph}$ denotes the fidelity of the multi-qubit phase gate, with infidelity $1-F_{\rm ph}\sim 1.8\times N/\sqrt{C}$, as extracted from the numerical simulations of  [55](Fig. 3(b,c)). $\mathcal{E}_{\rm l-GPG}$ ($\mathcal{E}^{\rm R}_{\rm l-GPG}$) is the (reverse of) preparation mapping associated with $W$ ($W^{\dagger}$) in the presence of losses, as detailed in Eq. (23). Here, for simplicity, we optimize parameters solely for the preparation process and apply the same parameters to the reverse mapping without additional optimization. The fidelity measure $F_{\rm pro}(\mathcal{E}_{\overline{H}},\overline{\mathcal{H}})$ is computed based on process matrices, which necessitates the superoperator formalism for each mapping in Eq. (29). The detailed calculations are provided in Appendix F.
In Fig. 2, we plot the infidelity for: the implementation error of the Hadamard gate (orange), and the preparation error (blue) for the PI-11 code as a function of the cooperativity $C$. The simulation is performed in the presence of loss errors as described in Eq. (21). Following the approximation in Ref. [55] (Eq. (25)), the infidelity of the $\overline{H}$ implementation can be estimated as $1-F_{\rm pro}(\mathcal{E}_{\overline{H}},\overline{\mathcal{H}})\sim 21.78\times N/\sqrt{C}$, where we consider $18$ non-trivial GPG operations for the preparation step and its inverse, with $\theta=\pi/2$. Similarly, the infidelity for state preparation is estimated by $1-\langle+_{\rm pi11}|\rho|+_{\rm pi11}\rangle\sim 9.99\times N/\sqrt{C}$, where $9$ non-trivial GPG gates are considered, again with $\theta=\pi/2$. Our simulation results indicate a somewhat better performance with the infidelity scaling as $1-F_{\rm pro}(\mathcal{E}_{\overline{H}},\overline{\mathcal{H}})\sim 8.29\times N/C^{0.4985}$ and $1-\langle+_{\rm pi11}|\rho|+_{\rm pi11}\rangle\sim 2.80\times N/C^{0.4953}$, indicating lower infidelity compared to the approximation.
@@ VI Approximation of the super golden gate
The most efficient single-qubit gate set is given by representations of the binary icosahedral group $2I$ (which can be implemented with the Pollatsek-Ruskai code) and the super golden gate $\tau_{60}$ given in (6). This efficiency is measured in terms of gate counts. Here we give an approximate decomposition of $\tau_{60}$ which can be achieved by code-switching between bg-PI codes, the Pollatsek-Ruskai code that implements 2I, and PI codes that implement the transversal $T$ gate.
@@
VII On the fault-tolerance of our scheme
Spin system errors:- An operator $Q$ is fault-tolerant on the spin-system for QEC codes on system $A$ and $B$ that correct $t$ errors if for any multiqubit Pauli operator $P$ of weight $t$, there exists an operator $A$ with weight at most $t$ such that $QP=AQ$. Operationally, this means that when a weight $t$ error $P$ propagates across a quantum circuit $Q$, the error $P$ evolves in the Heisenberg picture to an error $A$ that does not have too many errors. This definition of fault-tolerance on the spin-system is consistent with the definition of fault tolerance proposed in Ref [65]. For us, the 7 qubit and 11 qubit PI codes correct a single error, and hence $t=1$.
Each constituent gate in our GPGs and conditional GPGs is spin fault-tolerant, which demonstrates the spin fault-tolerance of our protocol. These constituent gates are the rotation operators $R(\theta\hat{J}^{z}_{\Gamma})$ and $R(\theta\hat{J}^{y}_{\Gamma})$ and the mode displacement operator. Mode displacement operators are trivially spin fault-tolerant because they commute with all spin errors.
To demonstrate the spin-fault tolerance of the rotation operators, it suffices to focus our attention on $R(\theta\hat{J}^{z}_{\Gamma})$. The rotation operator $R(\theta\hat{J}^{z}_{\Gamma})$ is trivially spin fault-tolerant with respect to $Z$ errors because it commutes with $Z$ errors on the spin system. From Ref. [66], the rotation operator also obeys the commutation relations $R(\theta\hat{J}^{z}_{\Gamma})X_{\Gamma,k}=e^{i\theta Z_{\Gamma,k}}X_{\Gamma,k}R(\theta\hat{J}^{z}_{\Gamma})$ and $R(\theta\hat{J}^{z}_{\Gamma})Y_{\Gamma,k}=e^{i\theta Z_{\Gamma,k}}Y_{\Gamma,k}R(\theta\hat{J}^{z}_{\Gamma})$ for $X$ and $Y$ errors respectively. These commutation relations imply that both an $X$ and $Y$ error on the $k$th qubit evolves under the rotation operator into an error that remains localized on the $k$th qubit. Hence the rotation operator $R(\theta\hat{J}^{z}_{\Gamma})$ is indeed spin fault-tolerant.
Mode errors:- Loss or gain errors can occur on our bosonic mode, which we model using the lowering operator $\hat{a}$ and the raising operator $\hat{a}^{\dagger}$ respectively. We also consider phase errors on the bosonic mode, which are generated by the number operator $\hat{a}^{\dagger}\hat{a}$. We say that an operator $Q$ is fault-tolerant on the bosonic system, or Bose fault-tolerant, if for any operator $\bar{B}$ that is either a ladder operator or a number operator, there exists an operator $A$ with weight at most $t$ on the spin systems such that $Q\bar{B}=AQ$.
The rotation operator $R(\theta\hat{w}_{\Gamma})$ obeys the commutation relations $R(\theta\hat{w}_{\Gamma})\hat{a}=e^{-i\theta\hat{w}_{\Gamma}}\hat{a}R(\theta\hat{w}_{\Gamma})$ and $R(\theta\hat{w}_{\Gamma})\hat{a}^{\dagger}=e^{i\theta\hat{w}_{\Gamma}}\hat{a}^{\dagger}R(\theta\hat{w}_{\Gamma})$. This shows that ladder operator errors on the bosonic mode can propagate to the spin system as unitaries. Hence the rotation operator is not Bose fault-tolerant with respect to ladder operators, unless $\theta$ is very close to an integer multiple of $2\pi.$ On the other hand the rotation operator is trivially Bose fault-tolerant with respect to the number operator.
While the mode displacement operator does not commute with the ladder operators, it does not introduce additional errors to the spin system because ladder operator errors still propagate to ladder operator errors. Hence the displacement operator is Bose fault-tolerant with respect to ladder operators. The displacement operator is furthermore Bose fault-tolerant with respect to number operators because it propagates the number operator to a linear combination of the number operator and ladder operators. Hence the displacement operator is Bose fault-tolerant.
To mitigate the impact of the rotation gate not being Bose fault-tolerant in general, we should aim to perform both the displacement gate and the rotation gate as quickly and as accurately as possible, minimizing the number of gain or loss errors. On the other hand, our protocol is fault-tolerant with respect to bosonic phase errors.
VIII Conclusion
@@
We have presented a method for code switching between a stabilizer code and a permutation invariant code, and along the way have introduced a family of PI codes with a tunable set of transversal, non-Clifford gates. It is noteworthy that our measurement-free code-switching scheme is also applicable to transitions between two stabiliser codes, provided they satisfy the even-odd quantum code structure. In comparison to the measurement-based approach, our scheme eliminates the overhead associated with measurements and reduces the gate operation complexity by utilising collective gate implementations. Given the recent development on how quantum error correction on arbitrary permutation-invariant codes can be implemented using geometric phase gates [41], we expect it to be possible to extend our work to a fully fault-tolerant setting, using ideas from fault-tolerance theory [65, 67, 68, 69]. Additionally, a numerical study of the efficiency of our protocol for fault-tolerant quantum computation is an interesting line of future line of enquiry that is currently beyond the scope of our paper.
IX Acknowledgements
@@
Y.O. acknowledges support from EPSRC (Grant No. EP/W028115/1) and also the EPSRC funded QCI3 Hub under Grant No. EP/Z53318X/1. G.K.B. and Y.J. acknowledge support from the Australian Research Council Centre of Excellence for Engineered Quantum Systems (Grant No. CE 170100009).
@@ Appendix A AAB+ code
$T_{b}^{\otimes n}|1_{b,g}\rangle$ $=e^{ig\pi/b}|{{1}}_{b,g}\rangle.$ (37)
@@ Appendix C Transversal non-Clifford logical operations
Appendix D Construction of codes from linear programming
Let us consider the problem of constructing quantum codes encoding a single logical qubit that correct against a set of Kraus operators $\{K_{i}\}$ given the promise that the orthogonality of the Knill-Laflamme quantum error correction conditions are satisfied.
Now suppose that the codespace resides in a subspace of the space spanned by the orthonormal vectors $\{|c_{j}\rangle\}$, and we write the logical codewords as linear combinations of these orthonormal vectors. Specifically, we demand that the logical codewords are linear combinations of distinct basis states given by
$|0_{L}\rangle$ $=\sum_{j\ {\rm even}}\sqrt{x_{j}}|c_{j}\rangle$ (45)
$|1_{L}\rangle$ $=\sum_{j\ {\rm odd}}\sqrt{x_{j}}|c_{j}\rangle,$ (46)
where the coefficients $x_{j}$ are real and non-negative.
Then
Such a method has been described in [38], and we explain in the next section how to apply this idea explicitly to construct extended bg codes that correct errors of increasing distance, and hence ensure the scalability of the codes.
Appendix E $(b,g,m)$-PI codes
Here, we prove that $(b,g,m)$-PI codes have distance $d\geq 2t+1$ whenever $g,2b-g\geq 2t+1$ and when $m\geq t$. Clearly, when $g,|2b-g|\geq d$, the code has a bit-flip distance of $d$. Hence, for the code to have a distance of $d$, it suffices to to ascertain that the non-deformation conditions of the Knill-Laflamme conditions holds for the diagonal errors, that is
$\langle 0_{b,g,m}|P_{j}|0_{b,g,m}\rangle=\langle 1_{b,g,m}|P_{j}|1_{b,g,m}\rangle,$ (47)
for all $k=0,\dots,d-1$, where $P_{j}=Z^{\otimes j}\otimes I^{\otimes N-j}$. Now we consider a code with logical codewords
$|0_{b,g,m}\rangle$ $=\sum_{k=0}^{m}a_{k}|D^{N}_{2bk}\rangle$
$|1_{b,g,m}\rangle$ $=\sum_{k=0}^{m}b_{k}|D^{N}_{2bk+g}\rangle,$ (48)
for real coefficients $a_{k},b_{k}$, with the idea of solving for $a_{k}$ and $b_{k}$ using linear algebra techniques, akin to the idea in Refs [70, 38]. Then we can write the non-deformation conditions as
$\sum_{\begin{subarray}{c}0\leq k\leq m\\ \end{subarray}}a_{k}^{2}\langle D_{2kb}^{N}|P_{k}|D_{2kb}^{N}\rangle=\sum_{\begin{subarray}{c}0\leq k\leq m\\ \end{subarray}}b_{k}^{2}\langle D_{g+2kb}^{N}|P_{k}|D_{g+2kb}^{N}\rangle,$ (49)
which is a system of linear equations in the variables $a_{k}^{2}$ and $b_{k}^{2}$. The Dicke inner products $\langle D^{N}_{w}|P_{k}|D^{N}_{w}\rangle$ can be written as Krawtchouk polynomials [45, Lemma 6], namely
$\langle D^{N}_{w}|P_{k}|D^{N}_{w}\rangle=K^{N}_{w}(k)/\binom{N}{w},$ (50)
where
$K^{N}_{k}(z)=\sum_{j=0}^{z}\binom{z}{j}\binom{N-z}{k-j}(-1)^{j}$ (51)
is a binary Krawtchouk polynomial. In the language of generating functions,
$K^{N}_{k}(z)=[x^{k}](1-x)^{z}(1+x)^{N-z},$ (52)
where $[x^{k}]f(x)$ denotes the coefficient of $x^{k}$ of a polynomial $f(x)$. From the above generating function, we can also see that
$K^{N}_{k}(z)=[y^{N-k}](y-1)^{z}(y+1)^{N-z},$ (53)
and hence we have the symmetry relation
$K^{N}_{N-k}(z)$ $=[x^{N-k}](1-x)^{z}(1+x)^{N-z}$
$=[y^{N-k}](-1)^{z}(y-1)^{z}(y+1)^{N-z}$
$=(-1)^{z}[y^{k}](y-1)^{z}(y+1)^{N-z}$
$=(-1)^{z}K^{N}_{k}(z).$ (54)
We can again rewrite this as a system of homogeneous linear equations, $A{\bf x}=0$, where
$A$ $=\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}0\leq k\leq m\\ \end{subarray}}|j\rangle\langle 2k|\langle D_{2kb}^{N}|P_{j}|D_{2kb}^{N}\rangle$
$\quad+\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}0\leq k\leq m\\ \end{subarray}}|j\rangle\langle 2k+1|\langle D_{g+2kb}^{N}|P_{j}|D_{g+2kb}^{N}\rangle,$
$=\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}0\leq k\leq m\\ \end{subarray}}|j\rangle\langle 2k|K_{2kb}^{N}(j)/\binom{N}{2kb}$
$\quad+\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}0\leq k\leq m\\ \end{subarray}}|j\rangle\langle 2k+1|K_{g+2kb}^{N}(j)/\binom{N}{g+2kb},$ (55)
[…]
$a_{k}$ $=\sqrt{x_{2k}}\quad{k\ {\rm even}}$ (56)
$b_{k}$ $=\sqrt{-x_{2k+1}}\quad{k\ {\rm odd}}.$ (57)
We choose $m$ as the smallest integer for which (E) has a solution.
Next, let us simplify the linear program (E). Consider a vector ${\bf y}=(y_{0},y_{1},\dots,y_{2(m+1)-1})$ such that
$y_{k}=-y_{2m+1-k}$ (58)
for all $k=0,\dots,2m+1$. Then we can write ${\bf y}$ as a vector that depends on $(m+1)$ variables, in the sense that
${\bf y}=(y_{0},-y_{m},y_{1},-y_{m-1},\dots,y_{m},-y_{0}).$ (59)
For any vector ${\bf a}$ of the form
${\bf a}=(a_{0},a_{m},a_{1},a_{m-1},\dots,a_{m},a_{0}),$ (60)
it follows that
${\bf a}^{T}{\bf y}=$ $y_{0}a_{0}-y_{m}a_{m}$
$+y_{1}a_{1}-y_{m-1}a_{m}-1$
++$\dots$++ ⋯ +
$+y_{m}a_{m}-y_{0}a_{0}$
$=$ $0$ (61)
From (54) and $\binom{N}{g+2bk}=\binom{N}{2bk}$, every odd row of $A$ has the form of (62). Hence, every odd row is orthogonal to ${\bf y}$, and it suffices to find a ${\bf y}$ that is orthogonal to every even row in $A$.
For any vector ${\bf b}$ of the form
${\bf a}=(b_{0},-b_{m},b_{1},-b_{m-1},\dots,b_{m},-b_{0}),$ (62)
it follows that
${\bf b}^{T}{\bf y}=$ $y_{0}b_{0}+y_{m}b_{m}$
$+y_{1}b_{1}+y_{m-1}b_{m}-1$
++$\dots$++ ⋯ +
$+y_{m}b_{m}+y_{0}b_{0},$ (63)
and the constraint that ${\bf b}^{T}{\bf y}=0$ is then equivalent to showing that
$\sum_{k=0}^{m}y_{k}b_{k}=0.$ (64)
Hence, we consider a submatrix of $A$ given by
$B=\sum_{\begin{subarray}{c}0\leq j\leq d-1\\ j\ {\rm odd}\end{subarray}}\sum_{k=0}^{m}K^{N}_{2bk}(j)/\binom{N}{2bk}|j\rangle\langle k|.$ (65)
Then, any $(y_{0},\dots,y_{m})$ in the nullspace of $B$ gives a vector ${\bf y}=(y_{0},-y_{m},y_{1},-y_{m-1},\dots,y_{m},-y_{0})$ that is also in the nullspace of $A$. To construct a valid generalized $(b,g)$-PI code, it suffices to require that $(y_{0},\dots,y_{m})$ is a non-negative vector, that is, $(y_{0},\dots,y_{m})\geq 0$.
When $d=2t+1$ and $m=t$, the matrix $B$ has $t+1$ columns and $t$ rows. Then, $B$ is guaranteed to have a non-trivial nullspace. Numerically, in all of the examples we explored, we find that $B$ is a full-rank matrix, and its nullspace of $B$ always admits non-negative vectors. Namely, $B=({\bf e},\bar{B})$, where $\bar{B}$ is a square invertible matrix, and hence a valid $(y_{0},\dots,y_{m})$ is proportional to $\bar{B}^{-1}(y_{0},\dots,y_{m})$. Hence, by setting $m=t$, it is often possible to have generalized bg codes.
We like to prove that this generalized $(b,g)$-PI code has distance at least $12t+$ 1, thereby correcting $t$ errors, whenever (1) $g,(b-g)\geq 2t+1$ and (2) $m\geq t$. Note that condition (1) imposes the bit-flip distance, and hence it remains to show the non-deformation condition for the phase errors. From the above A-matrix type of argument, we just need to ascertain that the vector
$\gamma_{b,g,m}=(\gamma_{b,g,m,0}^{2},\dots,\gamma_{b,g,m,m}^{2})$ (66)
resides in the null space of the matrix $B$, that is, $B\gamma_{b,g,m}=0$. For this, we need to show that the following is true for all $x=1,\dots,t$:
$\sum_{k=0}^{m}\frac{K^{N}_{2bk}(2x-1)}{\binom{N}{2bk}}\binom{m}{k}\gamma_{b,g,m,k}^{2}=0.$ (67)
The system of linear equations can equivalently be written as
$\sum_{k=0}^{m}\binom{m}{k}K^{N}_{2x-1}(2bk)\gamma_{b,g,m,k}^{2}=0,$ (68)
which is
$\sum_{k=0}^{m}\binom{m}{k}K^{N}_{2x-1}(2bk)(N/2b-g/b)_{(m-k)}(N/2b)_{(k)}=0,$ (69)
It then remains to prove the following combinatorial lemma.
Lemma 1.
Let $b,g,m$ and $x$ be positive integers where $x\leq m$. Let $N=2bm+g$. Then
$S\coloneqq\sum_{k=0}^{m}\binom{m}{k}K^{N}_{2x-1}(2bk)(N/2b-g/b)_{(m-k)}(N/2b)_{(k)}=0.$
Proof.
The Krawtchouk polynomial $K^{N}_{2x-1}(2bk)$ can be written as a coefficient of a generating function, in the sense that
$K^{N}_{2x-1}(2bk)=[t^{2x-1}](1-t)^{2bk}(1+t)^{N-2bk}.$ (70)
Now let $\alpha=N/2b-g/b$ and $\beta=N/2b$. Note that $\alpha+\beta=2N/2b-g/b=(N-g)/b=2m$. Also, we can write
$\binom{m}{k}\alpha_{(m-k)}\beta_{(k)}$ $=m!\binom{\alpha}{m-k}\binom{\beta}{k}.$ (71)
Then we can write
$S$ $=m!\sum_{k=0}^{m}\binom{\alpha}{m-k}\binom{\beta}{k}[t^{2x-1}](1-t)^{2bk}(1+t)^{N-2bk}$
$=m![t^{2x-1}](1+t)^{N}\sum_{k=0}^{m}\binom{\alpha}{m-k}\binom{\beta}{k}((1-t)/(1+t))^{2bk}.$ (72)
Now
$\sum_{k=0}^{m}\binom{\alpha}{m-k}\binom{\beta}{k}y^{k}=[s^{m}](1+s)^{\alpha}(1+sy)^{\beta}.$ (73)
Identifying $y=(1-t)/(1+t)$, to prevent higher powers of $(1-t)/(1+t)$ to appear in the generating function when we make the substitution $y=(1-t)/(1+t)$, we have that whenever $x\leq m$,
$S$ $=m![t^{2x-1}s^{m}](1+t)^{N}(1+s)^{\alpha}(1+s((1-t)/(1+t))^{\beta}$
$=m![t^{2x-1}s^{m}]E(t,s)$ (74)
where
$E(t,s)=(1+s)^{\alpha}((1+t)^{2b}+s(1-t)^{2b})^{\beta}$ (75)
is a formal power series in the variables $t$ and $s$, which allows us to write
$E(t,s)=\sum_{u,v\geq 0}e_{u,v}t^{u}s^{v}.$ (76)
Now
$E(-t,1/s)$ $=(1+1/s)^{\alpha}((1-t)^{2b}+s^{-1}(1+t)^{2b})^{\beta}$
$=s^{-\alpha-\beta}(s+1)^{\alpha}(s(1-t)^{2b}+(1+t)^{2b})^{\beta}$
$=s^{-\alpha-\beta}E(t,s)$
$=s^{-2m}E(t,s).$ (77)
Using (77), it follows that
$\sum_{u^{\prime},v^{\prime}\geq 0}e_{u,v}(-t)^{u^{\prime}}s^{-v^{\prime}}=\sum_{u,v\geq 0}e_{u,v}t^{u}s^{v-2m}.$ (78)
We are interested in the scenario where $v-2m=-m$ and $-v^{\prime}=-m$, which is equivalent to $v^{\prime}=v=m$. Then we equate the coefficients of the left and right side of the above equation, focusing on the coefficient of $s^{-m}t^{u}$ for odd $u$. Then we get
$e_{u,m}=-e_{u,m}$ (79)
for odd $u$, which means that $e_{u,m}=0$ for odd $u.$ Recalling our constraint $x\leq m$, which means that we also need to have $u\leq 2m-1$, this proves the result. ∎
Since Lemma 1 is true, this shows that our $(b,g,m)$-PI code has the stipulated distance property.
Appendix F Fidelity measure for Hadamard gate process